TAB2 need help on word problem
1.Investing her bonus. Donna invested her $33,000 bonus and received a total of $970 in interest after one year. If part of the money returned 4% and the reminder 2.25% then how much did she invest at each rate?

2.Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipts were $824 and twice as many adults tickets as students tickets were sold, then how many of each were sold?
please show work

Answer by checkley77(12844)

.04X+.0225(33,000-X)=970
.04X+742.50-.0225X=970
.0175X=970-742.5
.0175X=227.5
X=227.5/.0175
X=13,000 INVESTED @ 4%.
33,000-13,000=20,000 INVESTED @ 2.25%.
PROOF:
.04*13,000+.0225*20,000=970
520+450=970=
970=970
-----------------------------------------
3*2X+2*X=824
6X+2X=824
8X=824
X=824/8
X=103 STUDENT TICKETS WERE SOLD.
2*103=206 ADULT TICKETS WERE SOLD.
PROOF
3*206+2*103=824
618+206=824
824=824